# On planar sections of the dodecahedron

### v1

@@ Abstract.
In the analysis of three-dimensional biological microstructures such as organoids, microscopy frequently yields two-dimensional optical sections without access to their orientation. Motivated by the question of whether such random planar sections determine the underlying three-dimensional structure, we investigate a discrete analogue in which the ambient structure is the vertex set of a Platonic solid and the observed data are congruence classes of planar intersections. For the regular dodecahedron with vertex set $V$, we define the planar statistic of a subset $X\subseteq V$ of vertices as the distribution of isometry types of inclusions $\Pi\cap X\subseteq V$, and ask whether this statistic determines $X$ up to planar congruence. We show that this is not the case: there exist two non-congruent $7$-element subsets with identical planar statistics.
@@ 1. Introduction
We refer to this distribution as the planar statistic of $X$, see Definition 1. Our primary question is whether two subsets with identical planar statistics must be congruent in $\mathbb{R}^{3}$. By analogy to results of Mallows and Clark [4] on intercept distributions in the plane, non-uniqueness is expected. We establish such non-uniqueness in the case of the regular dodecahedron, see Theorem 3.

### v2

@@ Abstract.
In the analysis of three-dimensional biological microstructures such as organoids, microscopy frequently yields two-dimensional optical sections without access to their orientation. Motivated by the question of whether such random planar sections determine the underlying three-dimensional structure, we investigate a discrete analogue in which the ambient structure is the vertex set of a Platonic solid and the observed data are congruence classes of planar intersections. For the regular dodecahedron with vertex set $V$, we define the planar statistic of a subset $X\subseteq V$ of vertices as the distribution of isometry types of inclusions $\Pi\cap X\subseteq V$, and ask whether this statistic determines $X$ up to isometry. We show that this is not the case: there exist two non-isometric $7$-element subsets with identical planar statistics.
@@ 1. Introduction
We refer to this distribution as the planar statistic of $X$, see Definition 1. Our primary question is whether two subsets with identical planar statistics must be isometric in $\mathbb{R}^{3}$. By analogy to results of Mallows and Clark [4] on intercept distributions in the plane, non-uniqueness is expected. We establish such non-uniqueness in the case of the regular dodecahedron, see Theorem 3.